Circular No. 6687

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 970616 K. Hurley, Space Sciences Laboratory, University of California at Ber-

keley (UCB); C. Kouveliotou, Universities Space Research Association at Marshall Space Flight Center; and F. Marshall, Goddard Space Flight Center, on behalf of large interplanetary network (IPN), BATSE and XTE teams, report a refined error box for this source, based on triangulation of the almost-final Ulysses and BATSE data, as well as XTE data reported on IAUC 6683. The IPN annulus, determined from the Ulysses-BATSE time difference, has a radius of 49^{o}.813, is centered at α = 22^{h}10^{m}30^{s}, δ = −27^{o}25′.0 (equinox 2000.0) and has a full width of 2′.2. The intersection of this annulus with the XTE error region has corners at α = 1^{h}19^{m}36^{s}, δ = −5^{o}42′.0; 1^{h}19^{m}36^{s}, −5^{o}46′.3; 1^{h}18^{m}23^{s}, −5^{o}12′.4 and 1^{h}18^{m}17^{s}, −5^{o}14′.4. The optical and x-ray sources described below lie within this error box.

T. Murakami, R. Fujimoto, Y. Ueda and R. Shibata, Institute of Space and Astronautical Science, Kanagawa, report, on behalf of a large international collaboration: “We observed the XTE/IPN error box of GRB 970616 (IAUC 6683) with ASCA during June 20.15–21.44 UT for a net total of 50 ks. Four x-ray sources are detected in the ASCA SIS field of view at the following positions: α = 1^{h}18^{m}48^{s}, δ = −5^{o}26′.3; 1^{h}18^{m}42^{s}, −5^{o}27′.1; 1^{h}18^{m}39^{s}, −5^{o}18′.3 and 1^{h}18^{m}35^{s}, −5^{o}19′.8 (equinox 2000.0, error radius 1′). The first and the fourth source error regions are entirely inside the refined IPN annulus reported above; the third source is slightly and the second is entirely outside the annulus. Their average SIS countrates at 0.7– 7 keV are 3.0, 3.3, 1.6 and 1.6 counts/ks, respectively (1 count/ks roughly corresponding to 2.3 × 10^{−14} erg cm^{−2} s^{−1} at 2–7 keV for a Crab-like spectrum). The first source appears to be variable, its countrate decreasing by a factor of two from the first to the second half of the observation.”

T. Galama, University of Amsterdam (UoA); K. Hurley, T. Broadhurst and B. Frye, UCB; and J. van Paradijs, UoA and University of Alabama at Huntsville, report, on behalf of a large international collaboration: “We have made R-band images of the error box of GRB 970616 with Keck II on June 21.61, 23.59 and 24.61 UT. We find a variable object 3′′.2 east and 3′′.6 north of a relatively bright object located at α = 1^{h}18^{m}55^{s}.7, δ = −5^{o}27′23′′ (equinox 2000.0; position inferred from the Digital Sky Survey). A preliminary differential analysis shows that between the first and third observations the object faded by 1.5±0.2 mag to mag ∼ 24. This variable is inside the refined IPN/XTE error box above but lies ∼ 2′ from the best position of the variable ASCA source.”

1997 June 25  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Brian G. Marsden